# E-Calib: A Fast, Robust and Accurate Calibration Toolbox for Event Cameras

### v1

@@ V-C Pose Estimation Results
While in most cases the calibration reprojection error can serve as a good benchmark for our method, in some occasions calibration can converge to a local minimum with the reprojection error being relatively low but the obtained sensor parameters are not optimal. Thus, we also validate the proposed method in terms of the estimated extrinsics. The estimated camera pose by our method is compared to the sensor ground truth pose given by the OptiTrack motion capture system in terms of the mean absolute error. The results are reported in Table III on ECam_ACircles dataset in terms of the mean absolute translation $\hat{\eta}_{t}$ and rotation errors $\hat{\eta}_{r}$, and their corresponding standard deviations $\sigma_{t}$ and $\sigma_{r}$. Fig. 13 also shows the positioning plots compared against the OptiTrack for $davis\_3x9\_gdlight$.
Accurate pose estimation is directly correlated with robust tracking of the calibration targets centers. If ST-DBSCAN cluster centers are utilized as circle centers, $\zeta{r}$ grows to 0.44 pixels and $\hat{\eta}_{t}$ rises to 6.71 cm on $davis\_3x9\_gdlight$. This demonstrates the need for eRWLS refinement since ST-DBSCAN cluster centers are affected by severe sensor noise, especially in challenging lighting conditions. In addition, our method provides accurate estimates of the sensor pose and maintains a pose estimation error in poorly lighted environments of the same order of magnitude compared to data sequences collected at good lighting conditions. This shows robustness against the evvent camera noise, which increases with decreasing illumination.
@@ V-D Benchmarks
To evaluate the impact of our event-driven calibration tool, we compare our results to state-of-the-art methods that include E2VID network by Muglikar et al. [17], FireNet by Scheerlinck et al. [32], and the MR approach by Reinbacher et al. [44]. The methods are compared to the work of this paper in terms of the detection success rate, reprojection error, and pose error if calibration was successful by a designated method. Fig. 14 compares the detection performance of our method against the aforementioned works. An important point is that detection success rate was not evaluated on the work of Huang et al. [19] since the required number of pattern detections is preset before running calibration. We have set this value to 700 sufficient to perform proper calibration, to compare the obtained results fairly and such that the total number of pattern detections are within the same order of magnitude. Table IV lists our results compared to those works on the $davis\_3x9\_gdlight$ dataset. In addition, only the radial distortion were computed for the benchmarks, while tangential distortion was ignored for all methods since it is not computed by Huang et al. [19]. It is also worth mentioning that these approaches were tested on the $davis\_3x9\_gdlight$ by the codes publicly available by the authors of these papers.
@@
In addition to the intrinsic parameters, the obtained extrinsics were also validated against the ground truth pose of the camera and compared to the camera pose estimation results by our method, Huang’s method, E2VID, and FireNet. Fig. 15 shows the positioning plot on $davis\_3x9\_gdlight$ along the y-axis and Table V quantifies these results with the same metrics reported in Table III. Table V also reports the positioning results of E2VID and FireNet, where a large positioning error was obtained due to the sub-optimal intrinsic parameters. On the other hand, Huang’s method and our approach provide satisfactory results in terms of a mean positioning error. However, our calibration tool shows a smoother estimate of camera trajectory, demonstrating the capability of our method to provide more optimal sensor intrinsics. This occured due to the sub-optimal optimization of the lens distortion coefficients and illustrates that sub-pixel localization accuracy of the circle centers is necessary for optimal sensor calibration, which is attained by our method unlike Huang’s approach.
TABLE V: Mean absolute pose error by our method and the works of Muglikar et al. [17], Scheerlinck et al. [32], and Huang et al. [19] compared against ground truth camera pose obtained from OptiTrack.
@@
The average execution time of the proposed work for datasets on the DAVIS346 and DVXplorer is also reported for each algorithm block, see Table VI. Compared to E2VID and FireNet networks, which run offline and require GPU acceleration, our algorithm outperforms these approaches in execution time without the need for heavy computational resources. On the other hand, while the MR approach achieves real-time performance, it suffers in terms of feature extraction performance and detection success rate. Nevertheless, our approach can surely be utilized for online calibration of the sensor. In addition, real-time performance is unnecessary for intrinsic calibration as redundant frames can be captured in close time steps. Nevertheless, it is worth mentioning that our algorithm is written on Python and MATLAB and was executed on Ryzen 7 6800H (3.2 GHz) processor. Thus, our execution time can be further improved if implemented using C++, especially if execution in real-time is needed.
TABLE VI: Execution times for the system blocks.

### v2

@@ V-C Pose Estimation Results
While in most cases the calibration reprojection error can serve as a good benchmark for our method, in some occasions calibration can converge to a local minimum with the reprojection error being relatively low but the obtained sensor parameters are not optimal. Thus, we also validate the proposed method in terms of the estimated extrinsics. The estimated camera pose by our method is compared to the sensor ground truth pose given by the OptiTrack motion capture system in terms of the mean absolute error. The results are reported in Table IV on ECam_ACircles dataset in terms of the mean absolute translation $\hat{\eta}_{t}$ and rotation errors $\hat{\eta}_{r}$, and their corresponding standard deviations $\sigma_{t}$ and $\sigma_{r}$. Fig. 13 also shows the positioning plots compared against the OptiTrack for davis_3x9_cone_gdlight. In addition, the mean pose error on all DAVIS and DVXplorer data sequences for all calibration trajectories is reported in table VII. Accurate pose estimation is directly correlated with robust tracking of the calibration targets centers. If ST-DBSCAN cluster centers are utilized as circle centers, $\zeta_{r}$ grows to 0.44 pixels and $\hat{\eta}_{t}$ rises to 6.71 cm on davis_3x9_cone_gdlight. This demonstrates the need for eRWLS refinement since ST-DBSCAN cluster centers are affected by severe sensor noise, especially in challenging lighting conditions. Moreover, OLS, MLESAC [45], and Hough Transform with RANSAC [46] provide mean translation errors of 2.431 cm, 1.983, 1.912 cm on DAVIS data sequences, respectively, while our method provides a mean translation error of 0.973 cm. Our proposed method provides accurate estimates of the sensor pose and maintains a pose estimation error in poorly lighted environments of the same order of magnitude compared to data sequences collected at good lighting conditions. This shows robustness against the event camera noise, which increases with decreasing illumination.
@@ V-D Benchmarks
To evaluate the impact of our event-driven calibration tool, we compare our results to state-of-the-art methods that include E2VID network by Muglikar et al. [17], FireNet by Scheerlinck et al. [32], and the MR approach by Reinbacher et al. [48]. The methods are compared to the work of this paper in terms of the detection success rate, reprojection error, and pose error if calibration was successful by a designated method. Fig. 14 compares the detection performance of our method against the aforementioned works. An important point is that detection success rate was not evaluated on the work of Huang et al. [19] since the required number of pattern detections is preset before running calibration. We have set this value to 700 sufficient to perform proper calibration, to compare the obtained results fairly and such that the total number of pattern detections are within the same order of magnitude. Table VIII lists our results compared to those works on the davis_3x9_cone_gdlight dataset. In addition, only the radial distortion were computed for the benchmarks, while tangential distortion was ignored for all methods since it is not computed by Huang et al. [19]. It is also worth mentioning that these approaches were tested on the davis_3x9_cone_gdlight by the codes publicly available by the authors of these papers.
@@
In addition to the intrinsic parameters, the obtained extrinsics were also validated against the ground truth pose of the camera and compared to the camera pose estimation results by our method, Huang’s method, E2VID, and FireNet. Fig. 15 shows the positioning plot on davis_3x9_gdlight along the y-axis and Table IX quantifies these results with the same metrics reported in Table IV. Table IX also reports the positioning results of E2VID and FireNet, where a large positioning error was obtained due to the sub-optimal intrinsic parameters. On the other hand, Huang’s method and our approach provide satisfactory results in terms of a mean positioning error. However, our calibration tool shows a smoother estimate of camera trajectory, demonstrating the capability of our method to provide more optimal sensor intrinsics. This occured due to the sub-optimal optimization of the lens distortion coefficients and illustrates that sub-pixel localization accuracy of the circle centers is necessary for optimal sensor calibration, which is attained by our method unlike Huang’s approach.
TABLE IX: Mean absolute pose error by our method and the works of Muglikar et al. [17], Scheerlinck et al. [32], and Huang et al. [19] compared against ground truth camera pose obtained from OptiTrack.
@@
The average execution time of the proposed work for datasets on the DAVIS346 and DVXplorer is also reported for each algorithm block, see Table X. Compared to E2VID and FireNet networks, which run offline and require GPU acceleration, our algorithm outperforms these approaches in execution time without the need for heavy computational resources. On the other hand, while the MR approach achieves real-time performance, it suffers in terms of feature extraction performance and detection success rate. Nevertheless, our approach can surely be utilized for online calibration of the sensor. In addition, real-time performance is unnecessary for intrinsic calibration as redundant frames can be captured in close time steps. Nevertheless, it is worth mentioning that our algorithm is written in Python and MATLAB and was executed on Ryzen 7 6800H (3.2 GHz) processor. Thus, our execution time can be further improved if implemented using C++, especially if execution in real-time is needed. In addition, since we employ an exponential fitting function in eRWLS, the execution time can be further reduced by employing a polynomial approximation instead of the exponential function.
TABLE X: Execution times for the system blocks.